OpenStudy (anonymous):

4x2 = 12x + 40 can you please show me how to do this?

OpenStudy (saifoo.khan):

Best way, 4x2 = 12x + 40 4x^2 - 12x - 40 = 0 Get rid of 4 by taking commons, 4(x^2 -3x - 10) = 0 Divide 4 on both sides, you are left with....

OpenStudy (saifoo.khan):

(x^2 -3x - 10) = 0

OpenStudy (anonymous):

can i factor this?

OpenStudy (saifoo.khan):

Now factor, x^2 - 3x -10 =0 x^2 -5x + 2x - 10 = 0 take commons, x(x-5) +2(x-5) = 0 (x+2)(x-5) = 0

OpenStudy (saifoo.khan):

Finally, solve them for Zero, x = -2 x = 5

OpenStudy (anonymous):

thanks =D

OpenStudy (saifoo.khan):

No problem!! :)

OpenStudy (saifoo.khan):

Please hit >>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>^^^^^^^^^^^^

OpenStudy (anonymous):

hit the tofu for good luck!

OpenStudy (saifoo.khan):

LOL.